OpenStudy (anonymous):

A value of \theta between 0 and \pi that satisfies the equation sin^3(theta)-cos^3(theta) =1 Also if anyone knows a good site form me to improve my trigonometry I'd appreciate it.

OpenStudy (blockcolder):

\[\sin^3\theta-\cos^3\theta=(\sin\theta-\cos\theta)(\sin^2\theta+\sin\theta\cos\theta+\cos^2\theta)\\ =(\sin\theta-\cos\theta)(1+\sin\theta\cos\theta)\] and I'm pretty sure there's a sum-to-product for this one.

OpenStudy (anonymous):

Thanks man, I was putting it into wolfram but couldn't work out what I should be aiming for. pi/2 is the answer. sing(pi/2)-cos(pi/2)(sin(pi/2)cos(pi/2)+1=1 =(1-0)(1*0+1)=1 1=1 qed. Thanks again.

OpenStudy (blockcolder):

No prob. :)